OpenStudy (anonymous):

Change this standard form equation into slope intercept form, then name the slope and the y-intercept. 2x - 3y = 12

OpenStudy (helpkay):

ok so you know y =mx+b

OpenStudy (helpkay):

@thizznlilb are you there and are you in algebra 1

OpenStudy (anonymous):

@helpkay yes

OpenStudy (helpkay):

sorry i'm back so are you doing flvs algebra 1

OpenStudy (anonymous):

yeah @helpkay

OpenStudy (helpkay):

ok what lesson is this so i can refresh

OpenStudy (helpkay):

ok never mind i got something else

OpenStudy (helpkay):

so what I s the first thing you think you need to do @thizznlilb remember y=mx+b

OpenStudy (mathmale):

You want to rewrite 2x - 3y = 12 in the form y=mx+b. To do this, divide all three of the given terms by -3. What do y ou get?

OpenStudy (helpkay):

@mathmale she is not online

OpenStudy (helpkay):

or he